# Uncovering thermodynamic origin of counterflow and coflow instabilities in miscible binary superfluids

### v1

@@ IV.2 Nonlinear stage for counterflow and coflow instability
Figure 7: Nonlinear time of counterflow instability at $T/T_{c}=0.677$, $\nu=-0.2$. Initial superfluid velocity is $v_{y}=0.628$. Plotted region is [0,40]$\times$[0,40]. From top to bottom, plotted values are $|\mathcal{O}_{1}|/|\mathcal{O}_{1}|_{\mathrm{max}}$, $\theta_{1}$, $|\mathcal{O}_{2}|/|\mathcal{O}_{2}|_{\mathrm{max}}$ and $\theta_{2}$. As time goes by, initial small perturbations grows exponentially. Then a dark soliton forms and vortex nucleation occurs. Due to strong dissipation within the system, vortices annihilate with anti-vortices. During this process, the kinetic energy of superfluids is also dissipated and the final state is homogeneous binary superfluids with velocity below critical velocity $v_{c1}$. Plotted region is [0,40]$\times$[0,40].
@@
Figure 8: Nonlinear evolution of coflow instability at $T/T_{c}=0.677$, $\nu=-0.2$. Initial superfluid velocity is $v_{y}=1.257$. The same values and regions are plotted as those for counterflow case. Phenomenon is similar to that of counterflow case, except coflow instability develops much more slowly and mildly. The final state is homogeneous binary superfluids with velocity below critical velocity $v^{\prime}_{c1}$.
@@
From linear analysis we have uncovered that miscible binary superfluids with superflow velocity generically present dynamical instabilities that follow from thermodynamic instabilities. In this section we explore the nonlinear stage of counterflow and coflow instabilities by evolving the system forward in time using the full set of nonlinear EoMs. Explicit form of EoMs and time evolution scheme can be found in Appendix B.
@@
In Figure 7 and Figure 8, we show typical examples of such nonlinear time evolution for counterflow and coflow instabilities, respectively. Phenomena for both cases looks similar, except that the coflow instability develops much more slowly and mildly. Since the system is dynamically unstable, small perturbations initially grow exponentially over time and eventually drive the system into a highly inhomogeneous state. Then nonlinear dynamics takes over and a dark soliton forms. Due to the instability of the dark soliton, vortex nucleation begins to occur. As the vortex number grows, the system is set into full quantum turbulence. In the next stage, due to strong dissipation in holographic superfluids, vortices and anti-vortices annihilate with each other and the total vortex/anti-vortex number gradually decreases. During this process, the kinetic energy of superfluids is also dissipated and the overall superfluid velocity is lowered. After a long time, when vortices are all annihilated, the final state of system is again a homogeneous binary superfluid, but with a superfluid velocity below the critical velocity, which is $v_{c1}$ for the counterflow case and $v^{\prime}_{c1}$ for the coflow case. Such slowing down mechanism might be universal for dissipative superfluids. In single-component superfluids, the superfluid velocity is also lowered to below the Landau critical velocity by vortex nucleation and annihilation Lan et al. .
@@ V Universal scaling law of critical velocity
In the holographic theory, we have uncovered four critical velocities: $v_{c1}$, $v_{c2}$ for counterflow case and $v_{c1}^{\prime}$, $v_{c2}^{\prime}$ for coflow case. Interestingly, we find three of the four critical velocities scale as $\nu^{1/2}$, see Figure 9. And this scaling behavior persists for all temperatures. The only velocity which does not follow this scaling behavior is the critical velocity $v^{\prime}_{c2}$. $v^{\prime}_{c2}$ is nonzero even when $\nu$ vanishes and barely changes as $\nu$ decreases. This can be understood as that $v^{\prime}_{c2}$ corresponds to Landau critical velocity which is due to the whole superfluid moving relative to the normal component and thus persists even the two components do not interact.
Such similar scaling behavior between the two models prompts us to identify $\nu$ with $1-\Delta^{2}$. In the recent work about interface instability of immiscible binary superfluids by some of us An et al. (2024a), similar scaling behavior for interface width between the two models was found for small $\nu$ by identifying $\nu$ with $1-\Delta$. In fact, the two identifications are equivalent for small $\nu$. If we take $\nu=1-\Delta^{2}$, for small $\nu$ we would have
@@
The coefficient 1/2 has no effect on scaling laws. It is nontrivial to find the same identification and scaling law between weakly interacting and strongly interacting binary superfluids independently in dynamical instability analysis of miscible binary superfluids and stationary state of immiscible binary superfluids. This result suggests that the scaling law for critical velocity might be universal, no matter whether the superfluid is weakly interacting or strongly interacting.
Figure 9: Two critical velocities $v_{c1}$, $v_{c2}$ for counterflow case and the first critical velocity $v^{\prime}_{c1}$ for coflow case versus coupling strength $\nu$ at $T/T_{c}=0.677$. All critical velocities scale as $\nu^{1/2}$. This result also holds at other temperatures.
@@ VI Conclusion and discussion
We have studied the counterflow and coflow instabilities in miscible binary superfluids. The former one is general in two-component fluid systems, while the latter instability is only presented in non-Galilean invariant systems, such as dissipative systems. We first developed a hydrodynamic framework for binary superfluids in Section II. Based in the local thermodynamic stability and positivity of entropy production, we have shown that binary superfluids become linearly dynamically unstable whenever the matrix of static susceptibility $\chi_{AB}$ is no longer positive definite.
@@
We have demonstrated that the dynamical counterflow instability found from GPEs is essentially captured by our criterion and thus this instability is of thermodynamic origin. In the framework of GPE, there is no coflow instability in an isolated uniform system because of Galilean invariance. In contrast to GPE, the holographic superfluids describes strongly coupled superfluids at finite temperature and dissipation. We found rich dynamics from the holographic miscible binary superfluids.
@@
From linear analysis, we have found that for both counterflow and coflow cases, dynamical instability only occurs above some critical superfluid velocity, see e.g. Figures 3 and 5. It first occurs in sound mode for the counterflow case, while in dissipative mode for coflow case. As the superfluid velocity further increases, another critical velocity shows for each case. For the counterflow case, the presence of the additional critical velocity is due to order competing mechanism between the two components. Above the second critical velocity, free energy of the counterflow binary superfluid phase is higher than that of single component phase, therefore the binary superfluid phase becomes less thermodynamically favored. In accordance, the unstable mode above the second critical velocity becomes the gapped mode, which is unstable even at zero wave number, signaling a global instability. For coflow case, the second critical velocity corresponds to Landau critical velocity in single component superfluids. Above this critical velocity, one of the sound modes also becomes unstable. Nevertheless, the unstable dissipative mode always dominates the instability.
@@
By formulating hydrodynamics of binary superfluids, we have found that except $v_{c2}$, which is due to order competing, other three instabilities originate in the divergences of susceptibility matrix $\chi_{AB}$. In this sense, all these instabilities are thermodynamic origin.
@@
Then we have explored the nonlinear stage of these dynamical instabilities by evolving the systems forward in time. We found the phenomena for two kinds of instabilities are similar, see Figures 7 and 8. In linear stage, small perturbations grow exponentially and drive the system into inhomogeneous state and then nonlinearity takes over. Firstly, dark soliton forms. Due to instability of dark soliton, vortex nucleation occurs. Then overall velocity of the system is reduced due to vortex annihilation and dissipation. The final state of the evolution is a homogeneous binary supefluid with velocity below critical velocity. Such mechanism seems to be general in dissipative superfluids.
@@
We have limited ourselves to the superfluids with identical components under the simplifying assumption of colinearity. Nevertheless, there should be no conceptual obstacle generalising to general cases. Moreover, it is desirable to understand the range of application of the hydrodynamic framework. In particular, one could expand the hydrodynamic expansion to higher orders to compare with the non-linear stage of the system. In following work, we shall also explore properties of vortex in strongly interacting miscible binary superfluids and compare them with those of weakly interacting binary superfluids Ishino et al. (2013a, b); Kuopanportti et al. (2019). While we have uncovered the thermodynamic origin of counterflow and coflow instabilities in miscible binary superfluids, it would be interesting to generalize the discussion to the immiscible binary superfluids, for which one has to deal with a phase-separated configuration. The tension from the interface could play an important role for the instability of the system Armas et al. (2017).
@@ Acknowledgements.
This work was partly supported by the National Natural Science Foundation of China Grants No. 12075298, No. 12122513, No. 11991052 and No. 12047503. We thank Jilin University for hospitality during the School of Frontiers in Hydrodynamics, Effective Field Theory and Holographic Duality. We acknowledge the use of the High Performance Cluster at Institute of Theoretical Physics, Chinese Academy of Sciences.
Appendix A Holographic setup

### v2

@@ IV.2 Nonlinear time evolution of the counterflow and coflow instabilities
Figure 7: Nonlinear evolution of the counterflow instability at $T/T_{c}=0.677$, $\nu=-0.2$. The initial superfluid velocity is $v_{y}=0.628$. The region plotted is $\times[0,40]$. From top to bottom, the values plotted are $|\mathcal{O}_{1}|/|\mathcal{O}_{1}|_{\mathrm{max}}$, $\theta_{1}$, $|\mathcal{O}_{2}|/|\mathcal{O}_{2}|_{\mathrm{max}}$ and $\theta_{2}$. As time goes by, the initial small perturbations grows exponentially. Then a dark soliton forms and vortex nucleation occurs. Due to strong dissipation within the system, vortices annihilate with anti-vortices. During this process, the kinetic energy of the superfluid components is also dissipated and the final state is a homogeneous binary superfluid with a velocity below the critical velocity $v_{c1}$. The region plotted is $\times[0,40]$.
@@
Figure 8: Nonlinear evolution of the coflow instability at $T/T_{c}=0.677$, $\nu=-0.2$. The initial superfluid velocity is $v_{y}=1.257$. The same values and regions are plotted as those for the counterflow case. This dynamics is similar to that of counterflow case, except that the coflow instability develops much more slowly and mildly. The final state is a homogeneous binary superfluid with a velocity below the critical velocity $v^{\prime}_{c1}$.
@@
From our linear analysis, we showed that miscible binary superfluids with superflow velocity generically present dynamical instabilities that follow from thermodynamic instabilities. In this section we explore the nonlinear stage of the counterflow and coflow instabilities by evolving the system forward in time using the full set of nonlinear EoMs. The explicit form of the EoMs and time evolution scheme can be found in Appendix D.
@@
In Figure 7 and Figure 8, we show typical examples of such nonlinear time evolution for the counterflow and coflow instabilities, respectively. The dynamics for both cases looks similar, except that the coflow instability develops much more slowly and mildly. Since the system is dynamically unstable, small inhomogeneous perturbations initially grow exponentially over time and eventually drive the system into a highly inhomogeneous state. Then nonlinear dynamics takes over and a dark soliton forms. Due to the instability of the dark soliton, vortex nucleation begins to occur. [^17] Vortices only nucleate if the initial perturbation breaks translations in both directions. As the number of vortices increases, the system transitions into a quantum turbulent stage. In the next stage, due to strong dissipation in holographic superfluids, vortices and anti-vortices annihilate with each other and the total number of vortex/anti-vortex pairs gradually decreases. During this process, the superfluid kinetic energy is also dissipated and the overall superfluid velocity is lowered. After a long time, when all vortices are annihilated, the final state of system is again a homogeneous binary superfluid, but with a superfluid velocity below the critical velocity, respectively below $v_{c1}$ for the counterflow case and $v^{\prime}_{c1}$ for the coflow case. Such slowing down mechanism might be universal for dissipative superfluids. In single-component superfluids, the superfluid velocity is also lowered to below the Landau critical velocity by vortex nucleation and annihilation Lan et al. .
@@ V Universal scaling law of critical velocity
In the holographic theory, we have found four critical velocities: $v_{c1}$, $v_{c2}$ for the counterflow case and $v_{c1}^{\prime}$, $v_{c2}^{\prime}$ for the coflow case. Interestingly, we find that three of the four critical velocities scale as $\nu^{1/2}$, see Figure 9. This scaling behavior persists for all temperatures. The only velocity which does not follow this scaling behavior is the critical velocity $v^{\prime}_{c2}$. $v^{\prime}_{c2}$ is nonzero even when $\nu$ vanishes and changes little as $\nu$ decreases. This is because $v^{\prime}_{c2}$ corresponds to the Landau critical velocity, due to the superfluid moving as a whole relative to the normal component, and thus persists even when the two components do not interact.
Such similar scaling behavior between the two models prompts us to identify $\nu$ with $1-\Delta^{2}$. In the recent work on interface instability of immiscible binary superfluids by some of us An et al. (2024a), similar scaling behavior for the interface width between the two models was found for small $\nu$ by identifying $\nu$ with $1-\Delta$. In fact, the two identifications are equivalent for small $\nu$. Taking $\nu=1-\Delta^{2}$, for small $\nu$ we have
@@
The coefficient $1/2$ has no effect on the scaling laws. It is nontrivial to find the same identification and scaling law between weakly and strongly interacting binary superfluids independently in the analysis of dynamical instabilities of miscible binary superfluids and in the stationary state of immiscible binary superfluids. This result suggests that the scaling law for critical velocity might be universal, whether the superfluid is weakly or strongly interacting.
Figure 9: Two critical velocities $v_{c1}$, $v_{c2}$ for counterflow case and the first critical velocity $v^{\prime}_{c1}$ for the coflow case versus coupling strength $\nu$ at $T/T_{c}=0.677$. All critical velocities scale as $\nu^{1/2}$. This result also holds at other temperatures.
@@ VI Conclusion and discussion
We have studied the counterflow and coflow instabilities in miscible binary superfluids. The former is generic in two-component fluid systems, while the latter is only present in non-Galilean invariant systems, such as dissipative systems. We first developed a hydrodynamic framework for binary superfluids in Section II.1. Based on the local thermodynamic stability and positivity of entropy production, we have shown that the hydrodynamic theory of binary superfluids becomes linearly dynamically unstable whenever the matrix of static susceptibility $\chi_{AB}$ is no longer positive definite.
@@
We have demonstrated that the dynamical counterflow instability found from GPEs is reproduced by our criterion, which provides a macroscopic origin for it. In the framework of GPE, there is no coflow instability in an isolated uniform system because of Galilean invariance. In contrast to GPE, the holographic superfluids describe strongly-coupled, dissipative superfluids at finite temperature. We found holographic miscible binary superfluids feature rich dynamics.
@@
Linear analysis reveals that for both the counterflow and coflow cases, a dynamical instability only occurs above some critical superfluid velocity, see e.g. Figures 3 and 5. For the counterflow case, it first arises as a sound mode crosses to the upper half complex frequency plane, while for the coflow case it is a diffusive mode. As the superfluid velocity is further increased, another critical velocity shows for each case. For the counterflow case, this is due to phase competition. Above the second critical velocity, free energy of the counterflow binary superfluid phase is higher than that of single component phase, therefore the binary superfluid phase becomes less thermodynamically favored. In accordance, the unstable mode above the second critical velocity becomes the gapped mode, which is unstable even at zero wave number, signaling a global instability. For the coflow case, the second critical velocity corresponds to the Landau critical velocity in single component superfluids. Above this critical velocity, one of the sound modes also becomes unstable. Nevertheless, the unstable diffusive mode always dominates the instability.
@@
By formulating hydrodynamics of binary superfluids, we have found that except $v_{c2}$, which is due to order competing, the other three instabilities originate in the divergences of susceptibility matrix $\chi_{AB}$.
@@
Then we have explored the nonlinear stage of these dynamical instabilities by evolving the systems forward in time. We found the phenomena for two kinds of instabilities are similar, see Figures 7 and 8. In linear stage, small perturbations grow exponentially and drive the system into inhomogeneous state and then nonlinearity takes over. Firstly, a dark soliton forms. Due to the instability of dark soliton, vortices are nucleated. Then the overall velocity of the system is reduced due to vortex annihilation and dissipation. The final state of the evolution is a homogeneous binary superfluid with a velocity below the critical velocity. Such mechanism seems to be general in dissipative superfluids.
@@
We have limited ourselves to superfluids with identical components under the simplifying assumption of colinearity. Nevertheless, there should be no conceptual obstacle relaxing this assumption. Moreover, it is desirable to understand the range of application of the hydrodynamic framework. In particular, one could expand the hydrodynamic expansion to higher orders to compare with the non-linear stage of the system. In future work, we shall also explore the properties of vortices in strongly interacting miscible binary superfluids (see An and Li (2025)) and compare them with those of weakly interacting binary superfluids Ishino et al. (2013a, b); Kuopanportti et al. (2019). While we have uncovered the macroscopic origin of counterflow and coflow instabilities in miscible binary superfluids, it would be interesting to generalize the discussion to immiscible binary superfluids, for which one has to deal with a phase-separated configuration. The tension from the interface could play an important role for the instability of the system Armas et al. (2017).
Finally, it would be interesting to investigate whether phonon-phonon Beliaev scattering at zero temperature produces a linearly dynamical instability with the GPE when the critical velocity is crossed. This would be the zero temperature counterpart of the finite temperature dynamical instabilities we found using dissipative hydrodynamics and gauge/gravity duality.
@@ Acknowledgements.
B. G. thanks Laurent Sanchez-Palencia for helpful comments on the draft. This work was supported by the National Natural Science Foundation of China Grants No. 12525503, No. 12075298, No. 12122513, and No. 12447101. We thank Jilin University for hospitality during the School of Frontiers in Hydrodynamics, Effective Field Theory and Holographic Duality. We acknowledge the use of the High Performance Cluster at Institute of Theoretical Physics, Chinese Academy of Sciences.
Appendix A Review of fluid hydrodynamics
Hydrodynamics Landau and Lifshitz (1987); Chaikin and Lubensky (1995); Kovtun (2012); Forster (2018) posits that the collective dynamics of an interacting system on scales long compared to the local equilibrium scales (e.g. the mean collision time or the mean free path for a system composed of weakly-interacting particles) is captured by the conservation equations of the densities of conserved operators (charge $n$, energy $\epsilon$, momentum $g^{i}$…) following from the global symmetries of the system ($U(1)$, time and space translations…). These equations are generalized continuity equations, eg
$\partial_{t}n+\partial j^{i}=0$ (39)
for charge conservation, where $n(t,x^{i})$ is the charge density and $j^{i}(t,x^{i})$ the spatial charge current. The expectation values of the spatial fluxes (charge $j^{i}$ and energy $j_{\epsilon}^{i}$ currents, spatial stress tensor $\tau^{ij}$…) in the thermal ensemble can be expanded order by order in gradients of the densities, with each term featuring a so-called transport coefficient (the shear and bulk viscosities, thermal conductivity, etc.). For instance, the charge current is
$j^{i}=nv^{i}-\sigma_{0}\partial^{i}\mu-\alpha_{0}\partial^{i}T+O(\partial^{2})\,.$ (40)
$v^{i}$ is the fluid velocity and $\sigma_{0}$, $\alpha_{0}$ are electrothermal conductivities. [^18] In a Galilean system, we would have $\sigma_{0}=\alpha_{0}=0$ so that the charge current is full convective, but in a non-Galilean system, such as Graphene Crossno et al. (2016); Lucas et al. (2016); Lucas and Fong (2018) near the charge neutrality point, other processes can generate a charge current.
Together with the first law of thermodynamics $d\epsilon=Tds+\mu dn+v_{i}dg^{i}$, this allows to solve the continuity equations and to compute the hydrodynamic spectrum.
In thermodynamics, global equilibrium is characterized by a few macroscopic variables, such as the charge and the energy. The first law of thermodynamics $dE=TdS+\mu dN$ relates variations of the energy $E$, and charge number $N$ to variations of the entropy $S$ through state variables such as the temperature and the chemical potential $(T,\mu)$. The second law states that the entropy is maximized in global equilibrium; it never decreases in any spontaneous process $\Delta S\geq 0$. This means that the entropy is a Lyapunov function and the Lyapunov stability theorems LaSalle and Lefschetz (1961) guarantee that global equilibrium is stable: small perturbations away from it decay over time.
In its modern incarnation, hydrodynamics is viewed as the local, out-of-equilibrium extension of global thermodynamic equilibrium Chaikin and Lubensky (1995); Kovtun (2012). That is, it is postulated that at finite temperature and for an interacting system, there exists a local equilibration time and length over which the system can be coarse-grained and microscopic dynamics (of e.g. particles) are recast into the conservation equations of the densities of conserved operators (charge $n(t,x^{i})$, energy $\epsilon(t,x^{i})$, momentum $g^{i}(t,x^{i})$ densities…) following from the global symmetries of the system (U(1), time and space translations…). This hypothesis can be verified explicitly by a Boltzmann equation treatment of the dynamics of a weakly-interacting gas of particles, or in strongly-interacting systems with a gravitational dual by gauge-gravity duality (see Appendix B). The hydrodynamic regime is enhanced by the strength of interactions, since elementary constituents will equilibrate faster, and strongly-interacting systems exhibit a larger window of scales where hydrodynamics applies.
We thus assume that each ‘mesoscopic’ volume element of the system is in local thermodynamic equilibrium. An immediate consequence of this is that local versions of the first and second laws of thermodynamics can be formulated Nicolis (1979); Chaikin and Lubensky (1995)
$d\epsilon=Tds+\mu dn+v_{i}dg^{i}\,$, (41)
$T\partial_{t}s+T\partial_{i}j_{s}^{i}\equiv\Delta\geq 0\,$, (42)
where $T(t,x^{i})$, $\mu(t,x^{i})$, $v^{i}(t,x^{i})$ are the local temperature, chemical potential and fluid velocity, and $s(t,x^{i})$, $j_{s}^{i}(t,x^{i})$ the local entropy density and entropy current, respectively. We use the Einstein summation convention for which repeated indices should be summed, and Latin indices $i$ run over spatial coordinates.
The conservation equations take the form of generalized continuity equations
$\partial_{t}\langle O^{A}\rangle+\partial_{i}\langle j^{iA}\rangle=0\,$, (43)
for the set of local expectation values of the densities of conserved operators $O^{A}=\{\epsilon,n,g^{i}\}$. The angular brackets denote an expectation value $\langle O^{A}(t,x^{i})\rangle=Z^{-1}\textrm{Tr}\left[O^{A}(t,x^{i})e^{\beta H}\right]$ in a thermal ensemble with inverse temperature $\beta$ and Hamiltonian $H$. From now on, we drop the angular brackets for brevity. To be closed, the equations require constitutive relations for the spatial fluxes $j^{iA}(t,x^{i})=\{j^{i}_{\epsilon},j^{i}_{n},\tau^{ij}\}$ (energy current, charge current and spatial stress tensor). Writing these is in general a very complicated task but is made tremendously easier by the hypothesis of local thermodynamic equilibrium: as $j^{iA}$ are not conserved, their expectation values in local thermal equilibrium can be written in an expansion in gradients of $O^{A}$:
$j^{iA}=\sum_{n\geq 0}D^{AB}_{(n)i}O_{B}$ (44)
where $D^{AB}_{(n)i}$ is a differential operator of order $n$ weighted by the appropriate power of the local equilibration length. Simply put, we write $j^{iA}$ as a series order by order in gradients. Each subsequent derivative term suppressed compared to the lower orders so that spatial variations about local equilibrium are small (but not restricted to linear perturbations, as manifest in the Navier-Stokes equations). Each independent gradient term in the differential operator $D^{AB}_{(n)i}$ comes with an independent transport coefficient, and our task now becomes to classify and constrain them.
In simple systems, there are only a handful of such terms at first order in gradients. In Galilean fluids, they are the shear and bulk viscosities and the thermal diffusivity Landau and Lifshitz (1987); Chaikin and Lubensky (1995). They appear in the constitutive relations as:
$g^{i}=$ $mnv^{i}\,,\quad j_{s}^{i}=sv^{i}-\frac{\kappa}{T}\partial^{i}T\,$, (45)
$\tau^{ji}=$ $p\delta^{ij}+v^{i}g^{j}-\eta\left(\partial^{i}v^{j}+\partial^{j}v^{i}\right)$
$-\left(\frac{2}{d}\eta+\zeta\right)\delta^{ij}\partial\cdot v\,.$
Here $p=Ts+\mu n+v_{i}g^{i}-\epsilon$ is the pressure and $d$ the spatial dimension of the system.
In more complicated systems with a lower amount of symmetry, their number quickly proliferates. For instance, there are fourteen first-order coefficients in a $(3+1)$-dimensional relativistic (or Galilean) superfluid where the conservation of the particle number is spontaneously broken Bhattacharya et al. (2011). Classifying them implies enumerating all possible tensor structures, eliminating redundancies using the equations of motion (EoMs) at lower orders, and choosing a frame. The choice of frame amounts to a choice of an out-of-equilibrium definition for the temperature, chemical potential etc. Kovtun (2012), which are only unambiguously defined in equilibrium.
Once all of these are done, an essential step remains: the local version of the second law of thermodynamics (42) must be obeyed on all fluid configurations, which sets some combination of transport coefficients to zero and sets inequality constraints on others. Finally, the Onsager reciprocity relations can be imposed depending on the set of discrete symmetries of the system (such as invariance under parity or time reversal).
In Galilean fluids (45), the positivity of entropy production amounts to imposing that the quadratic form
$\begin{split}\Delta=&\zeta(\partial\cdot v)^{2}+\frac{\kappa}{T}(\partial_{i}T)^{2}\\ &+\eta\left(\partial^{i}v^{j}+\partial^{j}v^{i}-\frac{2}{d}\delta^{ij}\partial\cdot v\right)^{2}\end{split}$ (46)
is positive for all fluid configurations, resulting in the simple set of constraints $\eta,\zeta,\kappa\geq 0$.
Once the hydrodynamic theory is formulated, we Fourier transform in space and time, such that after linearizing around local equilibrium $O^{A}(t,x^{i})=O^{A}_{0}+\delta O^{A}e^{-i\omega t+ik_{i}x^{i}}$, the equations (43) take the form: [^19] The matrix $M$ only depends on $k^{i}$ and not $\omega$ provided the hydrodynamic frame, ie the out-of-equilibrium definition of temperature, chemical potential and fluid velocity Kovtun (2012), is chosen appropriately.
$(-i\omega+M(k^{i})\cdot\chi^{-1})\delta O=0\,$, (47)
where $\chi_{AB}\equiv\delta O_{A}/\delta s_{B}=\delta^{(2)}p/(\delta s_{A}\delta s_{B})$ is the matrix of static susceptibilities. Here $\delta s_{B}=(\delta T,T\delta(\mu/T),\delta v^{i})$ is the set of sources conjugate to $\delta O^{A}=(\delta\epsilon,\delta n,\delta g^{i})$. The static susceptibilities are just thermodynamic derivatives, for instance $\chi_{\epsilon\epsilon}=\left.\partial\epsilon/\partial T\right|_{(\mu/T)}$.
The spectrum of collective excitations $\omega(k)$ in the complex frequency plane ($k=\sqrt{k_{i}k^{i}}$) is determined by finding the $\omega(k)$ such that $\det(i\omega(k)-M\cdot\chi^{-1})=0$. Hydrodynamics only reliably captures poles such that $\omega(k\to 0)\to 0$, i.e. these poles can be expanded as $\omega(k)=\sum_{n>0}\omega_{n}k^{n}$. Since we expanded in plane waves $e^{-i\omega t+ik_{i}x^{i}}$, any any hydrodynamic pole in the upper half complex frequency plane $\textrm{Im}\,\omega>0$ leads to exponential growth of perturbations and signals the dynamical instability.
Returning to Galilean fluids (45), the hydrodynamic modes (thermal diffusion, transverse momentum diffusion and longitudinal sound) take relatively simple expressions Chaikin and Lubensky (1995) and it is straightforward to show by direct inspection of their closed form expressions that the constraints from positivity of entropy production, together with local thermodynamic stability (the matrix $\chi_{AB}$ is positive-definite) implies linear dynamical stability: all the hydrodynamic poles are in the lower complex frequency plane, $\textrm{Im}\,\omega(k)\leq 0$. On the other hand, if one of the eigenvalues of $\chi_{AB}$ changes sign, then a linear dynamical instability is found.
In a sense, this result is expected from the property that the entropy is a Lyapunov function and from the Lyapunov stability theorem. However, showing in practice that the modes are all linearly stable and that an instability occurs whenever $\chi_{AB}$ is no longer positive definite can be technically complicated by several factors: if many modes couple; the number of transport coefficients is not small; the system has a reduced amount of symmetry; or the frame choice is such that the linearized equations are not of the form (47) and spurious non-hydrodynamic unstable modes are present in the spectrum.
Appendix B Review of gauge/gravity duality
Gauge/gravity duality, also known as holography or AdS/CFT (Anti-de Sitter/Conformal Field Theory) correspondence, has been widely used to study strongly-correlated condensed matter systems without quasiparticles (see Zaanen et al. (2015) for a pedagogical textbook, Hartnoll et al. for more advanced applications to quantum matter and Ammon and Erdmenger (2015) for a more string-theoretic exposition). Gauge/gravity duality originates from string theory Maldacena (1998); Witten (1998); Gubser et al. (1998), which is a theory of quantum gravity, and states that certain conformal gauge field theories are dual to a classical gravity theory in the ‘t Hooft limit when the rank of the gauge group is taken to infinity and the coupling constant is large. This is both a strong/weak coupling and a quantum/classical duality, which makes it particularly appealing. The gravity theory is defined on asymptotically AdS spacetimes (the ‘bulk’). Anti de Sitter is a maximally symmetric solution to Einstein’s equations with a negative cosmological constant. It has a negative global Riemann curvature and so is hyperbolic, which endows it with a conformal timelike boundary Witten (1998). This is at the heart of this holographic duality as, loosely speaking, the energy scale of the dual field theory is geometrized as the radial coordinate interpolating between the center of the AdS spacetime and its boundary. The dual field theory has a lower dimensionality (that of the conformal boundary of the bulk spacetime) than the gravity theory. In this work, we consider only $(3+1)$-dimensional gravity theories, dual to $(2+1)$-dimensional field theories.
In practice, the conformal symmetry is broken, either by placing the theory at finite density or turning on various other deformations of the dual field theory. Finite temperature, finite density states in the dual field theory are modeled by black holes in the bulk charged under a global U(1) symmetry. They are mathematical solutions to the classical equations of the Einstein-Maxwell theory with a negative cosmological constant. For zero Maxwell field, the solution is the Schwarzschild black hole, while for nonzero Maxwell field it is the Reissner-Nordström black hole. The temperature in the dual field theory is given by the Hawking temperature of the black hole event horizon, while the charge is given through a generalized Gauss’s law by the electric flux due to the bulk Maxwell field coming from the black hole horizon.
Since the gauge theory is in the strong coupling regime, we use its correlation functions to characterize it. This is done by identifying the field theory generating functional $Z$ with the on-shell, renormalized, Euclidean gravity path integral according to the well-established holographic dictionary Hartnoll et al. ; Zaanen et al. (2015). Computing correlation functions on the field theory side amounts to solving the linearized Einstein and Maxwell equations and working out their responses to electromagnetic and gravitational perturbations. The bulk fields are ‘projected’ on the conformal boundary, and the sources and responses are read off from their asymptotic expansions. As an example, using $z$ to denote the radial coordinate interpolating between the boundary at $z=0$ and the horizon at $z=z_{h}$, the bulk Maxwell field is found to have the following near-boundary expansion:
$A_{\mu}(z,t,x,y)=a_{\mu}(t,x,y)+b_{\mu}(t,x,y)z+O(z^{2})\,.$ (48)
The indices $\mu=(z,t,x,y)$ run over the holographic radial coordinate $z$, the time coordinate $t$ and spatial coordinates $(x,y)$. The dual field theory has coordinates $(t,x,y)$. $a_{\mu}$ and $b_{\mu}$ are integration functions which are fixed by boundary conditions at $z=0$ and $z_{h}$. There are two since Maxwell’s equations are second-order PDEs. Subleading coefficients are entirely fixed by $a_{\mu}$ and $b_{\mu}$. $a_{t}$ is proportional to the chemical potential of the state, and $b_{t}$ to its charge density. More generally, the integration function leading in $z$ is the source (here, $a_{\mu}$), and the subleading one the response (here, $b_{\mu}$). Thus the current-current retarded Green’s functions is $G^{R}_{J^{\mu}J^{\nu}}\propto\delta b_{\mu}/\delta a^{\nu}$.
Since holographic systems are dual to strongly-coupled gauge theories, it is natural to expect that their late time, long distance collective dynamics are captured by hydrodynamics. This was indeed verified for systems dual to relativistic fluids first at linear level Policastro et al. (2001, 2002a, 2002b); Herzog (2003), and then at nonlinear level Bhattacharyya et al. (2008); Banerjee et al. (2011); Erdmenger et al. (2009). At linear level, the spectrum of collective hydrodynamic modes maps onto the quasinormal modes (QNMs) of the black hole perturbations Kovtun and Starinets (2005). These electromagnetic and gravitational waves (in asymptotically AdS spacetimes rather than Minkowsky as for astrophysical gravitational waves) are damped and dissipation is encoded by imposing that in the classical gravity approximation, they can only be absorbed by the black hole horizon Son and Starinets (2002).
How does one model a superfluid using Gauge-Gravity duality? This was answered in a series of papers Gubser (2008); Hartnoll et al. (2008b, a), where a charged, complex scalar field is added to the Einstein-Maxwell theory, which now has an extra classical Klein-Gordon equation for the complex scalar. This complex scalar field in the bulk is dual to a complex scalar operator in the boundary field theory, with a scaling dimension set by the mass of the bulk scalar. Below a certain critical temperature $T_{c}$, two solutions compete. One is the Reissner-Nordström black hole, with the scalar field vanishing everywhere in the bulk; the other has a non-vanishing bulk scalar field. Since the Klein-Gordon equation is second-order, the solution also comes with two free integration constants, one interpreted as the source of the scalar operator and the other as the scalar expectation value. Any nontrivial scalar profile breaks the global $U(1)$ symmetry explicitly for nonzero source. The insight in Gubser (2008); Hartnoll et al. (2008b, a) was that a solution can be found where the source is zero and the $U(1)$ symmetry is then broken spontaneously, giving rise to a superfluid phase in the dual field theory.
The match between the superfluid hydrodynamics and the collective thermal dynamics of these holographic superfluids was thoroughly established over a series of works Amado et al. (2009); Herzog and Yarom (2009); Sonner and Withers (2010); Herzog et al. (2011); Bhattacharya et al. (2011); Goutéraux and Mefford (2020a, b); Arean et al. (2021); Goutéraux et al. (2023); Areán et al. (2024).
Appendix C Holographic setup